### v1

# Towards a Design Guideline for RPA Evaluation: A Survey of Large Language Model-Based Role-Playing Agents

Abstract
Role-Playing Agent (RPA) is an increasingly popular type of LLM Agent that simulates human-like behaviors in a variety of tasks. However, evaluating RPAs is challenging due to diverse task requirements and agent designs. This paper proposes an evidence-based, actionable, and generalizable evaluation design guideline for LLM-based RPA by systematically reviewing 1,67616761,6761 , 676 papers published between Jan. 2021 and Dec. 2024. Our analysis identifies six agent attributes, seven task attributes, and seven evaluation metrics from existing literature. Based on these findings, we present an RPA evaluation design guideline to help researchers develop more systematic and consistent evaluation methods.
Towards a Design Guideline for RPA Evaluation:
A Survey of Large Language Model-Based Role-Playing Agents
Chaoran Chen University of Notre Dame                        Bingsheng Yao Northeastern University                        Ruishi Zou University of California, San Diego
Wenyue Hua University of California, Santa Barbara                        Weimin Lyu Stony Brook University                        Toby Jia-Jun Li University of Notre Dame
Dakuo Wang thanks: Corresponding author: d.wang@northeastern.edu      Equal contribution.    Github repository: https://github.com/CRChenND/LLM_roleplay_agent_eval_survey Northeastern University
Figure 1: RPA evaluation design guideline. To illustrate how to use it in practice, we pretended we were selecting the evaluation metrics for the “Stanford Agent Village” Park et al. (2023) given agent attributes (yellow) and task attributes (pink). The original authors’ selection of evaluation metrics (purple and blue) perfectly aligns with our RPA design guideline, which echoes their work’s robustness. More details in Sec 5.1 and a bad example in Sec 5.2.
1 Introduction
LLMs have yielded human-like performance in various cognitive tasks (e.g., memorization Schwarzschild et al. (2025), reasoning Wang et al. (2023); Plaat et al. (2024), and planning Song et al. (2023); Huang et al. (2024)). These emergent capabilities have fueled growing research interest on Role-Playing Agent (RPA) Chen et al. (2024d); Tseng et al. (2024): RPAs are digital intelligent agent systems powered by LLMs, where users provide human-like agent attributes (e.g., personas) and task attributes (e.g., task descriptions) as input, and prompt the LLM to generate human-like behaviors and the reasoning process. The potential of RPAs is promising and far-reaching, as illustrated by the early results of the massive interdisciplinary studies in social science Park et al. (2022, 2023), network science Chen et al. (2024b), psychologyJiang et al. (2024) and juridical science He et al. (2024b).
Despite growing interest in RPAs, a fundamental question remains: how can we systematically and consistently evaluate an RPA? How should we select the evaluation metrics, so that the evaluation results can be comparable or generalizable from one task to another task? Addressing these challenges is difficult Dai et al. (2024); Tu et al. (2024); Wang et al. (2024c). due to the vast diversity of tasks (e.g., simulating an individual’s online browser behavior Chen et al. (2024b) or simulating a hospital Li et al. (2024c)), and the high flexibility in RPA design (e.g., an agent persona can be one sentence or 2-hours of interview log Park et al. (2024)). Another challenge is the inconsistent and often arbitrary selection of evaluation methods and metrics for RPAs, raising concerns about the validity and reliability of evaluation results Wang et al. (2025b); Zhang et al. (2025). As a result, the research community finds it difficult to compare the performance across multiple RPAs in similar tasks reliably and systematically.
To address this gap, we propose an evidence-based, actionable, and generalizable design guideline for evaluating LLM-based RPAs. We conducted a systematic literature review of 1,67616761,6761 , 676 papers on the LLM Agent topic and identified $122122$ papers describing its evaluation details. Through expert coding, we found that agent attribute design interacts with task characteristics (e.g., simulating an individual or simulating a society requires a diverse set of agent attributes). Furthermore, we synthesized common patterns in how prior research successfully (or unsuccessfully) designed their evaluation metrics to correspond to the RPA’s agent attributes and task attributes. Building on these insights, we propose an RPA evaluation design guideline (Fig. 1) and illustrate its generalizability through two case studies.
2 Related Work
Existing surveys on the evaluation of RPAs Gao et al. (2024); Chen et al. (2024d); Tseng et al. (2024); Chen et al. (2024e); Mou et al. (2024a) provide a unified classification of RPA evaluation metrics from the perspective of evaluation approaches. However, they lack a comprehensive and consistent taxonomy for versatile evaluation metrics, leading to arbitrary metrics selection in practices.
Prior works Gao et al. (2024); Mou et al. (2024a) categorize RPA evaluations into three types: automatic evaluations, human-based evaluations, and LLM-based assessments. Automatic evaluations are efficient and objective, but lack context sensitivity, failing to capture nuances like persona consistency. Human-based evaluations provide deep insight into character alignment and engagement, but they are costly, less scalable, and prone to subjectivity. LLM-based evaluations are automatic and offer scalability and speed, but may not always align with human judgments.
The classification of evaluation metrics in prior works varies significantly, leading to inconsistency and ambiguity. For instance, Gao et al. (2024) focuses on realness validation and ethics evaluation, whereas Chen et al. (2024d) differentiates between character persona and individualized persona. Furthermore, Chen et al. (2024e) classifies evaluation into conversation ability, role-persona consistency, role-behavior consistency, and role-playing attractiveness, which partially overlap with Mou et al. (2024a)’s individual simulation and scenario evaluation. These discrepancies indicate a lack of standardized taxonomy, making it difficult to compare results across studies and select appropriate evaluation metrics for specific applications.
While existing surveys offer different taxonomies of RPA evaluation, they do not provide concrete evaluation design guidelines. Our work addresses this gap by proposing a structured framework that systematically links evaluation metrics to RPA attributes and real-world applications.
3 Method
We conduct a systematic literature review to address our research question. Following prior method Nightingale (2009), we aim to identify relevant research papers on RPAs and provide a comprehensive summary of the literature. We selected four widely used academic databases: Google Scholar, ACM Digital Library, IEEE Xplore, and ACL Anthology. These databases encompass a broad spectrum of research across AI, human-computer interaction, and computational linguistics. Given the rapid advancements in LLM research, we included both peer-reviewed and preprint studies (e.g., from arXiv) to capture the latest developments. Below, we detail our paper selection and annotation process.
Table 1: Definition and examples of six agent attributes.
Agent attributes Definition Examples
Activity History A record of past actions, behaviors, and engagements, including schedules, browsing history, and lifestyle choices. Backstory, plot, weekly schedule, browsing history, social media posts, lifestyle
Belief and Value The principles, attitudes, and ideological stances that shape an individual’s perspectives and decisions. Stances, beliefs, attitudes, values, political leaning, religion
Demographic Information Personal identifying details such as name, age, education, career, and location. Name, appearance, gender, age, date of birth, education, location, career, household income
Psychological Traits Characteristics related to personality, emotions, interests, and cognitive tendencies. Personality, hobby and interest, emotional
Skill and Expertise The knowledge level, proficiency, and capability in specific domains or technologies. Knowledge level, technology proficiency, skills
Social Relationships The nature and dynamics of interactions with others, including roles, connections, and communication styles. Parenting styles, interactions with players
3.1 Literature Search and Screening Method
Figure 2: Screening process of literature review. We initially retrieved 1,67616761,6761 , 676 papers published between 2021 and 2024, and narrowed down to 122122122122 final selections.
Our literature review focuses on LLM agents that role-play human behaviors, such as decision-making, reasoning, and deliberate actions. We specifically focus on studies where LLM agents demonstrate the ability to simulate human-like cognitive processes in their objectives, methodologies, or evaluation techniques. To ensure methodological rigor, we define explicit inclusion and exclusion criteria (Tab. 6 in Appendix A).
The inclusion criteria require that an LLM agent in the study exhibits human-like behavior, engages in cognitive activities such as decision-making or reasoning, and operates in an open-ended task environment. We excluded studies where LLM agents primarily serve as chatbots, task-specific assistants, evaluators, or agents operating within predefined and finite action spaces. Additionally, studies focusing solely on perception-based tasks (e.g., computer vision or sensor-based autonomous driving) without cognitive simulation were also excluded.
Using this scope, we searched four databases using the query string provided in Appendix B, retrieving 1,67616761,6761 , 676 papers published between January 2021 to December 2024. After removing duplicates, 1,57315731,5731 , 573 unique papers remained. Two authors independently screened the paper titles and abstracts based on the inclusion criteria. If at least one author deemed a paper relevant, it proceeded to full-text screening, where two authors reviewed the paper in detail and resolved any disagreements through discussion (Fig. 2). The final set of selected studies comprised $122122$ publications.
3.2 Paper Annotation Method
Our team followed established open coding procedures Brod et al. (2009) to conduct an inductive coding process to identify key themes. Three co-authors with extensive experience in LLM agents (“annotators,” hereinafter) collaboratively annotated the papers on three dimensions: agent attributes, task attributes, and evaluation metrics.
To ensure consistency, two annotators independently annotated the same 20% of articles and then held a meeting to discuss and refine an initial set of categories for the three dimensions. After reaching a consensus, each annotator annotated half of the remaining papers and cross-validated the other half annotated by the other annotator. Once the annotations were completed, a third annotator reviewed the coded data and identified potential discrepancies. Any discrepancies were discussed among the annotators to ensure consistency until disagreements were resolved, ensuring reliability and validity through an iterative refinement process.
4 Survey Findings
Building on the annotated data, we systematically categorized agent attributes, task attributes, and evaluation metrics. We then present a structured RPA evaluation design guideline, outlining how to select appropriate evaluation metrics based on agent and task attributes.
Table 2: Definition of seven task attributes.
Task attributes Definition
Simulated Individuals Simulating specific individuals or groups, such as users and participants.
Simulated Society Simulating social interactions, such as cooperation, competition, and communication.
Opinion Dynamics Simulating political views, legal perspectives, and social media content.
Decision Making Simulating decision-making of stakeholders in investment, public policies, or games.
Psychological Experiments Simulating human traits, including personality, ethics, emotions, and mental health.
Educational Training Simulating teachers and learners to enable personalized teaching and accommodate learner needs.
Writing Simulating readers or characters to support character development and audience understanding.
Table 3: Definitions and examples of seven evaluation metric categories.
Evaluation Metrics Definitions Examples
Performance Assess RPAs’ effectiveness in task execution and outcomes. Prediction accuracy
Psychological Measure human psychological responses to RPAs and the agents’ self-awareness and emotional state. Big Five Invertory
External Alignment Evaluate how closely RPAs align with external ground truth or human behavior and judgments. Alignment between model and human
Internal Consistency Assess coherence between an RPA’s predefined traits (e.g., personality), contextual expectations, and behavior. Personality-behavior alignment
Social and Decision-Making Analyze RPAs’ social interactions and decision-making, including their effects on negotiation, societal welfare, markets, and social dynamics. Social Conflict Count
Content and Textual Evaluate the quality, coherence, and diversity of RPAs’ text, including semantic understanding, linguistic style, and engagement. Content similarity
Bias, Fairness, and Ethics Assess biases, extreme or unbalanced content, or stereotyping behavior. Factual error rate
4.1 Agent Attributes
We identified six categories of agent attributes, as shown in Tab. 1. Activity history refers to an agent’s longitudinal behaviors, such as browsing history Chen et al. (2024b) or social media activity Navarro et al. (2024). Belief and value encompass the principles, attitudes, and ideological stances that shape an agent’s perspectives, including political leanings Mou et al. (2024c) or religious affiliations Lv et al. (2024). Demographic information includes personal details such as name, age, education, location, career status, and household income. Psychological traits include an agent’s personality Jiang et al. (2023a), emotions, and cognitive tendencies Castricato et al. (2024). Skill and expertise describe an agent’s knowledge and proficiency in specific domains, such as technology proficiency or specialized professional skills. Lastly, social relationships define the social interactions, roles, and communication styles between agents, including aspects like parenting styles Ye and Gao (2024) or relationships between players Ge et al. (2024).
4.2 Task Attributes
We identified seven key types of RPA downstream task attributes (Tab. 2). These tasks fall into two broad categories: those that use simulation as a research goal and those that use simulation as a tool to support specific research domains.
Among them, simulated individuals and simulated society primarily use simulation as the research goal. Simulated individuals involve modeling specific individuals or groups, such as end-users Chen et al. (2024a), to study their behaviors and interactions in a controlled setting. Simulated Society focuses on social interactions, including cooperation Bouzekri et al. (2024), competition Wu et al. (2024), and communication Mishra et al. (2023), aiming to explore emergent social dynamics.
In contrast, the other task attributes employ simulation as a means to serve specific research domains. Opinion dynamics entails simulating political views Neuberger et al. (2024), legal perspectives Chen et al. (2024c), and social media discourse Liu et al. (2024c) to analyze the formation and evolution of opinions. Decision making addresses the decision-making processes of stakeholders in investment Sreedhar and Chilton (2024) and public policy Ji et al. (2024), providing insights into strategic behaviors. Psychological experiments explore human traits such as personality Bose et al. (2024), ethics Lei et al. (2024), emotions Zhao et al. (2024), and mental health De Duro et al. (2025), using simulated scenarios to study cognitive and behavioral responses. Educational training supports personalized learning by simulating teachers and learners, enhancing pedagogical approaches and adaptive education systems Liu et al. (2024d). Finally, writing involves modeling readers or characters to facilitate character development Benharrak et al. (2024) and audience engagement Choi et al. (2024), contributing to storytelling and content generation research.
Figure 3: Proportional distribution of agent-oriented metrics across different agent attributes.
Figure 4: Proportional distribution of task-oriented metrics across different task attributes.
4.3 Agent- and Task-Oriented Metrics
We derived seven categories of evaluation metrics (Tab. 3) that are shared by agent- and task-oriented metrics despite differences in the specific metrics.
Agent Attributes Top 3 Agent-Oriented Metrics
Activity History External alignment metrics, internal consistency metrics, content and textual metrics
Belief and Value Psychological metrics, bias, fairness, and ethics metrics
Demographic Info. Psychological metrics, internal consistency metrics, external alignment metrics
Psychological Traits Psychological metrics, internal consistency metrics, content and textual metrics
Skill and Expertise External alignment metrics, internal consistency metrics, content and textual metrics
Social Relationship Psychological metrics, external alignment metrics, social and decision-making metrics
Table 4: Top 3 frequently used agent-oriented metrics for each agent attribute
Task Attributes Top 3 Task-Oriented Metrics
Simulated Individuals Psychological, performance, and internal consistency metrics
Simulated Society Social and decision-making metrics, performance metrics, and psychological metrics
Opinion Dynamics Performance metrics, external alignment metrics, and bias, fairness, and ethics metrics
Decision Making Social and decision-making, performance, and psychological metrics
Psychological Experiment Psychological, content and textual, and performance metrics
Educational Training Psychological, performance, and content and textual metrics
Writing Content and textual, psychological, and performance metrics
Table 5: Top 3 frequently used task-oriented metrics for each task attribute
Agent-oriented metrics focus on intrinsic, task-agnostic properties that define an RPA’s essential ability, such as underlying reasoning, consistency, and adaptability. These include performance metrics like memorization, psychological metrics such as emotional responses measured via entropy of valence and arousal, and social and decision-making metrics like social value orientation. Additionally, agent-oriented evaluations emphasize internal consistency metrics (e.g., consistency of information across interactions), external alignment metrics (e.g., hallucination detection), and content and textual metrics such as clarity. These evaluations ensure logical coherence, factual accuracy, and alignment with expected behavioral and cognitive frameworks, independent of any specific task.
Task-oriented metrics evaluate an RPA’s effectiveness in performing specific downstream tasks, focusing on task-related aspects such as accuracy, consistency, social impact, and ethical considerations. Performance measures how well RPAs execute designated tasks, such as prediction accuracy. Psychological metrics assess human psychological responses to RPAs, including self-awareness and emotional states; for example, the Big Five Inventory. External alignment evaluates how closely RPAs align with external ground truth or human behavior; for instance, alignment between model and human. Internal consistency ensures coherence between an RPA’s predefined traits, contextual expectations, and behavior; for example, personality-behavior alignment. Social and decision-making metrics analyze RPAs’ influence on negotiation, societal welfare, and social dynamics; for instance, the social conflict count. Content and textual quality focuses on the coherence, linguistic style, and engagement of RPAs’ generated text, such as content similarity. Lastly, bias, fairness, and ethics metrics examine biases, extreme content, or stereotypes; for instance, the factual error rate. Together, these seven metrics provide a comprehensive framework for evaluating RPAs’ task performance and broader impact.
4.4 RPA Evaluation Design Guideline
Building on our previous classification of agent attributes, task attributes, and evaluation metrics, we observed that both agent design and evaluation can be broadly divided into two categories: agent-oriented and task-oriented. This distinction led us to investigate patterns between agent design and evaluation, aiming to develop systematic guidelines for selecting evaluation metrics in future research.
Step 1. Selecting Agent-Oriented Metrics Based on Agent Attributes
We analyzed the distribution of agent attributes and agent-oriented metrics, as illustrated in Fig. 3. Our analysis reveals that, for each agent attribute, the top three categories of agent-oriented metrics account for the majority of all metric types. Based on this observation, our first guideline recommends selecting agent-oriented metrics according to agent attributes. Specifically, we suggest referring to Tab. 4 to identify the top three corresponding metrics. For instance, for Activity History, the recommended metrics are external alignment, internal consistency, and content and textual metrics. Likewise, for Beliefs and Values, the most relevant choices are psychological metrics and bias, fairness, and ethics metrics. In particular, there are no established agent-oriented evaluation metrics for social relationships. Based on Social Exchange Theory Cropanzano and Mitchell (2005), which explains relationship formation through reciprocal interactions and resource exchanges, we propose assessing social relationships with psychological metrics, external alignment metrics, and social and decision-making metrics.
Step 2: Selecting Task-Oriented Metrics Based on Task Attributes
Additionally, we analyzed the distribution of task attributes and task-oriented metrics, as shown in Fig. 4. Consistent with our previous findings, we observed that for each category of task attributes, the top three task-oriented metrics account for the vast majority of all metrics. Based on this, our second guideline recommends selecting task-oriented metrics according to task attributes. Specifically, we suggest referring to Tab. 5 to identify the top three corresponding metrics. For instance, for the Simulated Society task, the recommended metrics are social and decision-making, performance, and psychological metrics. Similarly, for the Opinion Dynamics task, the most relevant choices are performance, external alignment, bias, fairness, and ethics metrics.
However, these two steps should not be treated as one-time decisions. As the agent design process evolves, evaluation results may prompt adjustments to the attributes of the agent and the task, thereby influencing the selection of evaluation metrics. Therefore, this two-step evaluation guideline should be used iteratively to ensure that the evaluation remains adaptive to changing agent capabilities and task requirements. This iterative approach enhances the reliability, relevance, and robustness of RPA evaluation experiments.
5 Case Study: How to Use RPA Design Guideline to Select Evaluation Metrics
We present two case studies to illustrate how following our evaluation guidelines leads to the selection of a comprehensive set of evaluation metrics, while significant deviations may result in incomplete evaluation. By adopting the perspective of the original authors, we compare the evaluation outcomes resulting from adhering to or deviating from the RPA evaluation guidelines.
5.1 A Good Example: Generative Agents: Interactive Simulacra of Human Behavior
As shown in Fig. 1, Park et al. (2023) designed agents with demographic information, action history, and social relationships to create an interactive artificial society. Their evaluation methods are in line with the structured selection process proposed in our survey. Since no established agent-oriented evaluation metrics exist for social relationships, they focused on demographic information and action history. Referring to Fig. 3, they identified four relevant metric categories: Content and textual metrics, Internal consistency metrics, External alignment metrics, and Psychological metrics. Based on Tab. LABEL:tab:_long_table_agent_metrics_src in Appendix E, they selected five specific evaluation metrics: Self-knowledge (Content and textual, Internal consistency), Memory and Plans (Internal consistency), Reactions (External alignment), and Reflections (Psychological).
For task-oriented metrics, they determined that the agents’ downstream tasks aligned with simulated society and designed the evaluation metrics that are aligned with the top three most relevant metric types reported in Fig. 4. As shown in Tab. LABEL:tab:_long_table_task_metrics_src in Appendix E, they selected four evaluation metrics: Response accuracy (Performance), Relationship formation (Psychological), Information diffusion and Coordination (Social and decision-making). By systematically aligning evaluation metrics with agent attributes and task objectives, this approach ensured a comprehensive and meaningful assessment.
5.2 A Flawed Example: A Generative Social World for Embodied AI
A flawed example is presented in Appendix D Fig. 8, which is an ICLR submission, and the reviews are publicly available on OpenReview. The authors developed agents with demographic attributes, action history, psychological traits, and social relations for route planning and election campaigns. However, their evaluation deviated significantly from our RPA evaluation design guidelines.
Despite designing agents with clear attributes, they did not include any agent-oriented evaluation metrics. For task-oriented metrics, they identified tasks related to Opinion Dynamics and Decision-Making, which should have been evaluated using five key categories: Performance metrics, Psychological metrics, External alignment metrics, Social and decision-making metrics, and Bias, fairness, and ethics metrics. Instead, their evaluation relied solely on Arrival rate, Time, and Alignment between campaign strategies, leading to an incomplete assessment. This omission resulted in criticism from reviewers, as one noted: “The paper performs almost no quantitative experiments… This actually shows that the benchmark cannot cover too many current research methods, which is the biggest weakness of the paper.”
Figure 5: Relationships between agent attributes and downstream tasks. The numbers in the heatmap represent the paper counts.
6 Relationships Between Agent Attributes and Downstream Tasks
Both agent attributes and downstream task attributes play a crucial role in selecting appropriate RPA evaluation metrics. Researchers predefine these factors when designing and evaluating RPAs, yet their interrelation remains an open question. In this section, we analyze how agent attributes correspond to different downstream tasks, uncovering several recurring patterns (Fig. 5).
Demographic information and psychological traits are fundamental across all downstream tasks. Whether in decision-making, opinion dynamics, or simulated environments, these attributes consistently shape RPA design. As shown in Fig. 5, they are the most frequently incorporated factors, underscoring their central role in modeling agent behavior across diverse applications.
For tasks where simulation itself is the primary objective, such as Simulated Individuals and Simulated Society, the selection of agent attributes becomes broader. In addition to demographic and psychological factors, these tasks frequently incorporate skills, expertise, and social relationships, reflecting the need for richer agent representations to capture complex social and individual interactions. By contrast, tasks that use simulation as a means to study specific research fields tend to prioritize certain agent attributes. For instance, in Opinion Dynamics, beliefs and values play a distinctive role, as they directly influence how agents interact and form opinions. Similarly, tasks related to Educational Training and Writing exhibit a different pattern, emphasizing skills and expertise over broad demographic or psychological considerations.
In contrast, attributes such as activity history and social relationships receive significantly less emphasis across tasks. This raises a question: is their impact inherently limited, or are they simply underexplored in current RPA applications?
Overall, these findings highlight the nuanced interplay between agent attributes and downstream tasks. While demographic information and psychological traits are universally relevant, attributes like beliefs and values gain importance in specific contexts. At the same time, the relative absence of activity history and social relationships in current evaluations presents an open research question, particularly in scenarios requiring long-term modeling and complex social interactions.
7 Discussion
7.1 RPA: an Algorithm v.s. a System
Unlike traditional algorithmic innovations in NLP, the design of RPAs can not only support technical innovations to improve LLMs’ humanoid capabilities but also enable RPA-based simulation systems for practical benefits. For instance, from the perspective of psychology, RPAs support the exploration of human cognitive and behavioral activities in controlled yet highly scalable experiments, even in hypothetical scenarios. In social science, RPAs can deployed as proxies or pilot experiments to analyze and audit social systems, power dynamics, and human societal behaviors at scale. For the machine learning community, RPAs shed light on dynamic and human-centered model evaluations that are aligned with real-world scenarios by incorporating human and societal factors into consideration. Last but not least, HCI researchers are particularly intrigued by the implications of RPA systems that can provide personalized assistance with human-centered applications in various sectors, such as medicine, healthcare, and education.
Nevertheless, RPAs’ capability and flexibility are a double-edged sword; they not only have the potential to bring benefits to stakeholders but also expose potential risks and even harm if not responsibly designed. To what extent do RPAs’ responses align with genuine human cognitive activities, whether the cultural, linguistic, and contextual biases learned from the training data of LLMs impact predicted behaviors, and how to ensure RPAs’ robustness and consistency under different scenarios, are critical but under-explored challenges for both technical developers and system designers.
As a result, the design of RPAs should incorporate system design considerations while advancing technical explorations. For instance, RPA design should focus on target users from the very beginning of system design, emphasize the diversity of user backgrounds and perspectives, and iteratively refine the system, as suggested by Gould and Lewis (1985) and Shneiderman and Plaisant (2010) in established design guidelines for system usability. Nevertheless, differences in cultural norms, linguistic subtleties, and domain-specific knowledge can introduce variability in how RPAs are designed and perceived. Designers and developers must focus on a balance between generalization and specificity to ensure RPAs are both adaptable and effective across a wide range of scenarios.
7.2 The Design of RPA Persona
One of RPAs’ key strengths is their ability to adapt to diverse personas, tasks, and environments. But how can RPA personas be designed to ensure that LLMs faithfully and believably reflect the agents’ cognitive behaviors within a given task? Persona descriptions must strike a careful balance between intrinsic agent characteristics and contextual factors, ensuring thoughtful consideration of both the agents’ intrinsic characteristics and the contextual information of the specific environments for which the agents are designed.
The intrinsic characteristics of RPAs, such as their personal characteristics, education experience, domain expertise, emotional expressiveness, and decision-making processes, must be aligned with the purpose of the applications of RPAs. For example, an RPA designed for psychological experiments should prioritize cognitive characteristics like personality and empathy ability, whereas an RPA developed for economic simulations might emphasize negotiation tactics, competitive reasoning, and adaptability to changing conditions.
On the other hand, contextual information, such as task- and scenario-specific details, factors, and specifications, is equally critical in shaping the behaviors of RPAs. In healthcare applications, for instance, RPAs may simulate caregivers’ emotional responses to patients’ changing health status but still operate under clinical protocols, such as the ICU visitor rules. The granularity and fidelity of contextual information heavily influence the believability and effectiveness of the agents’ behaviors.
7.3 The Challenges of RPA Evaluation
The versatility of RPAs, which allows them to function in diverse roles and contexts, makes it infeasible to have a “one-solution-fits-all” evaluation metric for systematically evaluate RPAs both within and across tasks and user scenarios. One major difficulty lies in designing and determining task-oriented and agent-oriented evaluation metrics. Despite our work recommending an RPA evaluation design guideline based on a comprehensive review of the literature, existing evaluation metrics may not be sufficient to measure the performance of RPAs for different domain-specific applications.
The diversity of user scenarios further exacerbates the evaluation challenge. Different tasks may prioritize different aspects of RPAs, making it difficult to develop a one-size-fits-all evaluation framework. For instance, RPAs designed for psychological research focus on believable emotional responses, whereas RPAs for policymaking simulations underscore robustness to policy changes.
Moreover, cross-task evaluations pose significant challenges due to inconsistencies in how metrics are designed and applied across studies. The lack of standardized evaluation criteria complicates systematic benchmarking in RPA development and impedes interdisciplinary collaboration.
Addressing these challenges will require the development of systematic, multi-faceted evaluation frameworks that can accommodate the diverse applications and capabilities of RPAs while providing consistency and comparability across studies.
8 Conclusion
RPA evaluation lacks consistency due to varying tasks, domains, and agent attributes. Our systematic review of 1,67616761,6761 , 676 papers reveals that task-specific requirements shape agent attributes, while both task characteristics and agent design influence evaluation metrics. By identifying these interdependencies, we propose guidelines to enhance RPA assessment reliability, contributing to a more structured and systematic evaluation framework.
Limitations
RPAs are rapidly evolving and have widespread applications across various domains. While we aim to comprehensively review existing literature, we acknowledge certain limitations in our scope. First, our review may not encompass all variations of RPA evaluation approaches across different application domains. Second, new research published after December 2024 is not included in our analysis. As a result, our work does not claim to exhaustively cover all potential evaluation metrics. Instead, our goal is to provide a structured framework and actionable guidelines to help future researchers design more systematic and consistent RPA evaluations, even as the field continues to evolve.
Ethics Statement
Our work focuses on summarizing and analyzing the evaluation of RPAs, which we believe will be valuable to researchers in AI, HCI, and related fields such as psychological simulation, educational simulation, and economic simulation. We have taken care to ensure that this survey remains objective and balanced, neither overestimating nor underestimating trends. We do not anticipate any ethical concerns that arise from the research presented in this paper.
Appendix A Inclusion and Exclusion Criteria
We summarize the inclusion and exclusion criteria in Table 6. Briefly, the Inclusion Criteria (IC) ensure that the reviewed studies focus on LLM agents exhibiting human-like behavior—either implicitly (e.g., preference or behavioral patterns) or explicitly (e.g., emotions or personality)—along with key cognitive processes such as reasoning and decision-making. Moreover, an open-ended action space and the capacity to tackle multifaceted tasks are essential attributes for inclusion.
By contrast, the Exclusion Criteria (EC) eliminate studies employing LLMs purely as chatbots, single-purpose systems, or evaluation tools, rather than as agents mimicking human cognition. Likewise, if the LLM agents are restricted to fixed, close-ended tasks or limited to algorithmic optimization without simulating cognitive processes, they fall outside the scope of this work.
Table 6: Inclusion and exclusion criteria.
Inclusion Criteria (IC)
IC-1 The LLM agents in the paper simulate humanoid behavior with implicit personality (e.g., preference and behavior pattern) or explicit personality (e.g., emotion or characteristics).
IC-2 The LLM agents in the paper have cognitive activities such as decision-making, reasoning, and planning.
IC-3 The LLM agents in the paper are capable of completing complicated and general tasks.
IC-4 The LLM agents’ action set in the paper is neither predefined nor finite.
Exclusion Criteria (EC)
EC-1 The study does not employ LLM agents for simulation purposes but rather uses them as chatbots, task-specific agents, or evaluators.
EC-2 The paper’s research objectives, methodologies, and evaluations are not focused on simulating human-like behavior with LLM agents, but rather on optimizing LLM algorithms.
EC-3 The study primarily investigates the perception or action capabilities of LLM agents without simulating the cognitive process.
EC-4 The LLM agents are restricted to handling specific, close-ended tasks.
EC-5 The LLM agents’ actions are either predefined or limited.
Appendix B Query String
We employed the following query to guide our literature retrieval process:
(“large language model” OR LLM) AND (agent OR persona OR "human digital twin" OR simulacra) AND (simulat* OR generat* OR eval*) AND “human behavior” AND cognit*
This query was designed to capture a broad spectrum of studies on large language models that simulate or replicate human-like behavior. It combines keywords related to LLM agents (LLM, persona, simulacra), their capabilities (simulat*, generat*, eval*), and the focus on cognitively grounded human behavior (cognit*). This ensures that the resulting literature is relevant to our exploration of how LLM-based systems can mimic or exhibit human-like cognition and behavior patterns.
Appendix C Evaluation Approach Usage for Agent- and Task-Oriented Metrics
We present a breakdown of evaluation approach usage by agent-oriented metrics (Fig. 6) and task-oriented metrics (Fig. 7).
Figure 6: Usage ratio of evaluation approaches for each category of agent-oriented metrics.
Figure 7: Usage ratio of evaluation approaches for each category of task-oriented metrics.
Appendix D Case Study: Flawed Example
Fig. 8 visualized how the authors in the flawed example selected their evaluation metrics how further evaluation metrics could be uncovered through our proposed guideline.
Figure 8: Case study of a flawed example in Section 5.2. Given agent attributes (yellow) and task attributes (pink). The original authors’ selection of evaluation metrics (purple and blue). The missing metrics that are recommended by our proposed guideline (orange) align with the reviewer’s criticism in red text.
Appendix E Metrics Glossary
We present two glossary tables for referencing the source of agent-oriented metrics (Tab. LABEL:tab:_long_table_agent_metrics_src) and task-oriented metrics (Tab. LABEL:tab:_long_table_task_metrics_src).
Table 7: Agent-oriented evaluation metrics glossary.
Attribute Category Agent-oriented Metrics Approach Source
Belief & Value Bias, fairness, ethics metrics Exaggeration (normalized average cosine similarity) Automatic Cheng et al. (2023)
Belief & Value Bias, fairness, ethics metrics Individuation (classification accuracy) Automatic Cheng et al. (2023)
Belief & Value Bias, fairness, ethics metrics Bias (performance disparity, prevalence, magnitude, variation, attitude shift) Automatic Gupta et al. (2024)
Belief & Value Bias, fairness, ethics metrics Bias (performance disparity, prevalence, magnitude, variation, attitude shift) Automatic Taubenfeld et al. (2024)
Demographic Information Bias, fairness, ethics metrics Exaggeration (normalized average cosine similarity) Automatic Cheng et al. (2023)
Demographic Information Bias, fairness, ethics metrics Individuation (classification accuracy) Automatic Cheng et al. (2023)
Demographic Information Bias, fairness, ethics metrics Bias (performance disparity, prevalence, magnitude, variation, attitude shift) Automatic Gupta et al. (2024)
Demographic Information Bias, fairness, ethics metrics Bias (performance disparity, prevalence, magnitude, variation, attitude shift) Automatic Neuberger et al. (2024)
Demographic Information Bias, fairness, ethics metrics Bias (performance disparity, prevalence, magnitude, variation, attitude shift) Automatic Taubenfeld et al. (2024)
Demographic Information Bias, fairness, ethics metrics Message toxicity Automatic Fang et al. (2024)
Activity History Content and textual metrics Coherence LLM Li et al. (2024e)
Activity History Content and textual metrics Clarity Human Chen et al. (2024b)
Activity History Content and textual metrics Diversity of dialog (Shannon entropy, intra-remote-clique, inter-remote-clique, semantic similarity, longest common subsequence similarity) Automatic Ha et al. (2024)
Belief & Value Content and textual metrics Diversity of dialog (Shannon entropy, intra-remote-clique, inter-remote-clique, semantic similarity, longest common subsequence similarity) Automatic Gu et al. (2024)
Demographic Information Content and textual metrics Coherence LLM Li et al. (2024e)
Demographic Information Content and textual metrics Attitudes (topic term frequency) Automatic Fang et al. (2024)
Demographic Information Content and textual metrics Diversity of dialog (Shannon entropy, intra-remote-clique, inter-remote-clique, semantic similarity, longest common subsequence similarity) Automatic Fang et al. (2024)
Demographic Information Content and textual metrics Clarity Human Chen et al. (2024b)
Demographic Information Content and textual metrics Diversity of dialog (Shannon entropy, intra-remote-clique, inter-remote-clique, semantic similarity, longest common subsequence similarity) Automatic Ha et al. (2024)
Demographic Information Content and textual metrics Linguistic complexity (utterance length, Kolmogorov complexity) Automatic Milička et al. (2024)
Psychological Traits Content and textual metrics Text similarity (BLEU, ROUGE) Automatic Zeng et al. (2024)
Psychological Traits Content and textual metrics Tone Alignment LLM Zeng et al. (2024)
Skills and Expertise Content and textual metrics Coherence LLM Li et al. (2024e)
Activity History External alignment metrics Hallucination LLM Shao et al. (2023)
Activity History External alignment metrics Entailment LLM Li et al. (2024e)
Activity History External alignment metrics Believability/Credibility(self-knowledge, memory, plans, reactions, reflections) Human Park et al. (2023)
Demographic Information External alignment metrics Entailment LLM Li et al. (2024e)
Demographic Information External alignment metrics Believability/Credibility(self-knowledge, memory, plans, reactions, reflections) Human Park et al. (2023)
Psychological Traits External alignment metrics Fact Accuracy LLM Zeng et al. (2024)
Skills and Expertise External alignment metrics Hallucination LLM Shao et al. (2023)
Skills and Expertise External alignment metrics Entailment LLM Li et al. (2024e)
Activity History Internal consistency metrics Stability LLM Shao et al. (2023)
Activity History Internal consistency metrics Consistency of information Human Chen et al. (2024b)
Belief & Value Internal consistency metrics Attitude shift LLM Wang et al. (2024e)
Demographic Information Internal consistency metrics Stability LLM Shao et al. (2023)
Demographic Information Internal consistency metrics Attitude shift LLM Neuberger et al. (2024)
Demographic Information Internal consistency metrics Attitude shift LLM Taubenfeld et al. (2024)
Demographic Information Internal consistency metrics Behavior stability (mean, standard deviation) Automatic Wang et al. (2024g)
Demographic Information Internal consistency metrics Consistency of information Human Chen et al. (2024b)
Demographic Information Internal consistency metrics Consistency of psychological state / personalities Human Chen et al. (2024b)
Demographic Information Internal consistency metrics Consistency of information Human Zeng et al. (2024)
Psychological Traits Internal consistency metrics Stability LLM Shao et al. (2023)
Psychological Traits Internal consistency metrics Consistency of information Human Zeng et al. (2024)
Psychological Traits Internal consistency metrics Consistency of psychological state / personalities Human Zeng et al. (2024)
Psychological Traits Internal consistency metrics Consistency of information Human Cai et al. (2024)
Psychological Traits Internal consistency metrics Consistency of psychological state / personalities Human Cai et al. (2024)
Skills and Expertise Internal consistency metrics Stability LLM Shao et al. (2023)
Activity History Performance metrics Memorization LLM Shao et al. (2023)
Demographic Information Performance metrics Memorization LLM Chen et al. (2024b)
Demographic Information Performance metrics Communication ability (win rates) Automatic Liu et al. (2024a)
Demographic Information Performance metrics Reaction (accuracy) Automatic Liu et al. (2024a)
Demographic Information Performance metrics Self-knowledge (accuracy) Automatic Liu et al. (2024a)
Activity History Psychological metrics Empathy Human Chen et al. (2024b)
Belief & Value Psychological metrics Value LLM Shao et al. (2023)
Demographic Information Psychological metrics Personality consistency Automatic Wang et al. (2024c)
Demographic Information Psychological metrics Measured alignment for personality Human Wang et al. (2024c)
Demographic Information Psychological metrics Sentiment Automatic Fang et al. (2024)
Demographic Information Psychological metrics Empathy Human Chen et al. (2024b)
Demographic Information Psychological metrics Belief (stability, evolution, correlation with behavior) Automatic Lei et al. (2024)
Psychological Traits Psychological metrics Personality Automatic Shao et al. (2023)
Psychological Traits Psychological metrics Belief (stability, evolution, correlation with behavior) Automatic Shao et al. (2023)
Psychological Traits Psychological metrics Emotion responses (entropy of valence and arousal) Automatic Shao et al. (2023)
Psychological Traits Psychological metrics Personality (Machine Personality Inventory, PsychoBench) Automatic Jiang et al. (2023a)
Psychological Traits Psychological metrics Personality (vignette tests) Human Jiang et al. (2023a)
Belief & Value Social and decision-making metrics Social value orientation (SVO-based Value Rationality Measurement) Automatic Zhang et al. (2023b)
Table 8: Task-oriented evaluation metrics glossary.
Task Category Task-oriented Metrics Approach Source
Decision Making Social and economic metrics Negotiation (Concession Rate, Negotiation Success Rate, Average Negotiation Round) Automatic Huang and Hadfi (2024)
Decision Making Social and economic metrics Societal Satisfaction (average per-capita living area size, average waiting time, social welfare) Automatic Ji et al. (2024)
Decision Making Social and economic metrics Societal Fairness (variance in per capita living area size, number of inverse order pairs in house allocation, Gini coefficient) Automatic Ji et al. (2024)
Decision Making Social and economic metrics Macroeconomic (Inflation rate, Unemployment rate, Nominal GDP, Nominal GDP growth, Wage inflation, Real GDP growth, Expected monthly income, Consumption) Automatic Li et al. (2024d)
Decision Making Social and economic metrics Market and Consumer (Purchase probability, Expected competing product price, Customer counts, Price consistency between competitors) Automatic Gui and Toubia (2023)
Decision Making Social and economic metrics Market and Consumer (Purchase probability, Expected competing product price, Customer counts, Price consistency between competitors) Automatic Zhao et al. (2023)
Decision Making Social and economic metrics Probability weighting Automatic Jia et al. (2024)
Decision Making Social and economic metrics Utility (Intrinsic Utility, Joint Utility) Automatic Huang and Hadfi (2024)
Decision Making Psychological metrics Level of trust (distribution of amounts sent, trust rate) Automatic Xie et al. (2024a)
Decision Making Psychological metrics Risk preference Automatic Jia et al. (2024)
Decision Making Psychological metrics Loss aversion Automatic Jia et al. (2024)
Decision Making Psychological metrics Selfishness (Selfishness Index, Difference Index) Automatic Kim et al. (2024)
Decision Making Performance metrics Frequency (distribution of expert type) Automatic Wang et al. (2024b)
Decision Making Performance metrics Valid response rate Automatic Xie et al. (2024a)
Decision Making Performance metrics Web search quality (Mean reciprocal rank, Mean reciprocal rank) Automatic Ren et al. (2024a)
Decision Making Performance metrics Performance deviations/alignment from the baseline (accuracy, Jaccard Index, Cohen’s Kappa Coefficient, Percentage Agreement, overlapping ratio between prediction and targets) Automatic Kim et al. (2024)
Decision Making Performance metrics Performance deviations/alignment from the baseline (accuracy, Jaccard Index, Cohen’s Kappa Coefficient, Percentage Agreement, overlapping ratio between prediction and targets) Automatic Jin et al. (2024)
Decision Making Performance metrics Performance deviations/alignment from the baseline (accuracy, Jaccard Index, Cohen’s Kappa Coefficient, Percentage Agreement, overlapping ratio between prediction and targets) Automatic Wang et al. (2024b)
Decision Making Performance metrics Performance deviations/alignment from the baseline (accuracy, Jaccard Index, Cohen’s Kappa Coefficient, Percentage Agreement, overlapping ratio between prediction and targets) Automatic Wang et al. (2024f)
Decision Making Internal consistency metrics Behavioral alignment (lottery rate, behavior dynamic, Imitation and differentiation behavior, Proportion of similar and different dishes) Automatic Xie et al. (2024a)
Decision Making Internal consistency metrics Behavioral alignment (lottery rate, behavior dynamic, Imitation and differentiation behavior, Proportion of similar and different dishes) Automatic Zhao et al. (2023)
Decision Making Internal consistency metrics Cultural appropriateness (Alignment between persona information and its assigned nationality) LLM Li et al. (2024e)
Decision Making External alignment metrics Factual hallucinations (String matching overlap ratio) Automatic Wang et al. (2024f)
Decision Making External alignment metrics Simulation capability (Turing test) Human Ji et al. (2024)
Decision Making External alignment metrics Entailment LLM Li et al. (2024e)
Decision Making External alignment metrics Realism LLM Li et al. (2024e)
Educational Training Psychological metrics Perceived reflection on the development of essential non-cognitive skills Human Yan et al. (2024)
Educational Training Psychological metrics Non-cognitive skill scale Automatic Yan et al. (2024)
Educational Training Psychological metrics Sense of immersion / Perceived immersion Human Lee et al. (2023)
Educational Training Psychological metrics Perceived intelligence Human Cheng et al. (2024)
Educational Training Psychological metrics Perceived enjoyment Human Cheng et al. (2024)
Educational Training Psychological metrics Perceived trust Human Cheng et al. (2024)
Educational Training Psychological metrics Perceived sense of connection Human Cheng et al. (2024)
Educational Training Psychological metrics Personality (Big Five Invertory, MBTI score, SD3 score, Linguistic Inquiry and Word Count framework, HEXACO) Automatic Sonlu et al. (2024)
Educational Training Psychological metrics Personality (Big Five Invertory, MBTI score, SD3 score, Linguistic Inquiry and Word Count framework, HEXACO) Automatic Liu et al. (2024d)
Educational Training Psychological metrics Perceived usefulness Human Cheng et al. (2024)
Educational Training Performance metrics Density of knowledge-building Automatic Jin et al. (2023)
Educational Training Performance metrics Effectiveness of questioning Human Shi et al. (2023)
Educational Training Performance metrics Success criterion function outputs before operation and after operation Human Li et al. (2023a)
Educational Training External alignment metrics Knowledge level (reconfigurability, persistence, and adaptability) Automatic Jin et al. (2023)
Educational Training External alignment metrics Perceived human-likeness Human Cheng et al. (2024)
Educational Training Content and textual metrics Story Content Generation (narratives staging score) Automatic Yan et al. (2024)
Educational Training Content and textual metrics Willingness to speak Human Shi et al. (2023)
Educational Training Content and textual metrics Authenticity Human Lee et al. (2023)
Opinion Dynamics Psychological metrics Opinion change Human Triem and Ding (2024)
Opinion Dynamics Psychological metrics Emotional density Automatic Gao et al. (2023)
Opinion Dynamics Performance metrics Prediction accuracy (F1 score, AUC, MSE, MAE, depression risk prediction accuracy, suicide risk prediction accuracy) Automatic Gao et al. (2023)
Opinion Dynamics Performance metrics Prediction accuracy (F1 score, AUC, MSE, MAE, depression risk prediction accuracy, suicide risk prediction accuracy) Automatic Mou et al. (2024c)
Opinion Dynamics Performance metrics Prediction accuracy (F1 score, AUC, MSE, MAE, depression risk prediction accuracy, suicide risk prediction accuracy) Automatic Yu et al. (2024)
Opinion Dynamics Performance metrics Classification accuracy Human Chan et al. (2023)
Opinion Dynamics Performance metrics Rephrase accuracy Automatic Ju et al. (2024)
Opinion Dynamics Performance metrics Legal articles evaluation (precision, recall, F1) Automatic He et al. (2024a)
Opinion Dynamics Performance metrics Judgment evaluation for civil and administrative cases (precision, recall, F1) Automatic He et al. (2024a)
Opinion Dynamics Performance metrics Judgment evaluation for criminal cases (accuracy) Automatic He et al. (2024a)
Opinion Dynamics Performance metrics Prediction error rate Automatic Gao et al. (2023)
Opinion Dynamics Performance metrics Locality accuracy Automatic Ju et al. (2024)
Opinion Dynamics Performance metrics Decision probability Human Triem and Ding (2024)
Opinion Dynamics Performance metrics Decision volatility Human Triem and Ding (2024)
Opinion Dynamics Performance metrics Case complexity Human Triem and Ding (2024)
Opinion Dynamics Performance metrics Alignment (compare simulation results with actual social outcomes) Automatic Wang et al. (2024g)
Opinion Dynamics Internal consistency metrics Alignment (stance, content, behavior, static attitude distribution, time series of the average attitude) Automatic Mou et al. (2024c)
Opinion Dynamics Internal consistency metrics Personality-behavior alignment Human Navarro et al. (2024)
Opinion Dynamics Internal consistency metrics Similarity between initial and post preference (KL-divergence, RMSE) Automatic Namikoshi et al. (2024)
Opinion Dynamics Internal consistency metrics Role playing Human Lv et al. (2024)
Opinion Dynamics External alignment metrics Correctness Human He et al. (2024a)
Opinion Dynamics External alignment metrics Accuracy (correctness) Automatic Ju et al. (2024)
Opinion Dynamics External alignment metrics Logicality Human He et al. (2024a)
Opinion Dynamics External alignment metrics Concision Human He et al. (2024a)
Opinion Dynamics External alignment metrics Human likeness index Automatic Chuang et al. (2023b)
Opinion Dynamics External alignment metrics Alignment between model and human (Kappa correlation coefficient, MAE), Authenticity (alignment of ratings between the agent and human annotators) Human Chan et al. (2023)
Opinion Dynamics External alignment metrics Alignment between model and human (Kappa correlation coefficient, MAE), Authenticity (alignment of ratings between the agent and human annotators) Human Triem and Ding (2024)
Opinion Dynamics External alignment metrics Alignment between model and human (Kappa correlation coefficient, MAE), Authenticity (alignment of ratings between the agent and human annotators) Human Lv et al. (2024)
Opinion Dynamics Content and textual metrics Turn-level Kendall-Tau correlation (naturalness, coherence, engagingness and groundedness) Automatic Chan et al. (2023)
Opinion Dynamics Content and textual metrics Turn-level Spearman correlation (naturalness, coherence, engagingness and groundedness) Automatic Chan et al. (2023)
Opinion Dynamics Bias, fairness, and ethic metrics Partisan bias Automatic Chuang et al. (2023b)
Opinion Dynamics Bias, fairness, and ethic metrics Bias (cultural, linguistic, economic, demographic, ideological) Automatic Qu and Wang (2024)
Opinion Dynamics Bias, fairness, and ethic metrics Bias (mean) Automatic Chuang et al. (2023a)
Opinion Dynamics Bias, fairness, and ethic metrics Extreme values Automatic Chuang et al. (2023b)
Opinion Dynamics Bias, fairness, and ethic metrics Wisdom of Partisan Crowds effect Automatic Chuang et al. (2023b)
Opinion Dynamics Bias, fairness, and ethic metrics Opinion diversity Automatic Chuang et al. (2023a)
Psychological Experiment Social and economic metrics Money allocation Automatic Lei et al. (2024)
Psychological Experiment Psychological metrics Attitude change Automatic Wang et al. (2025a)
Psychological Experiment Psychological metrics Average happiness value per time step Automatic He and Zhang (2024)
Psychological Experiment Psychological metrics Belief value Automatic Lei et al. (2024)
Psychological Experiment Psychological metrics Personality (Big Five Invertory, MBTI score, SD3 score, Linguistic Inquiry and Word Count framework, HEXACO) Automatic He and Zhang (2024)
Psychological Experiment Psychological metrics Personality (Big Five Invertory, MBTI score, SD3 score, Linguistic Inquiry and Word Count framework, HEXACO) Automatic de Winter et al. (2024)
Psychological Experiment Psychological metrics Personality (Big Five Invertory, MBTI score, SD3 score, Linguistic Inquiry and Word Count framework, HEXACO) Automatic Bose et al. (2024)
Psychological Experiment Psychological metrics Personality (Big Five Invertory, MBTI score, SD3 score, Linguistic Inquiry and Word Count framework, HEXACO) Automatic Jiang et al. (2023b)
Psychological Experiment Psychological metrics Longitudinal trajectories of emotions Automatic De Duro et al. (2025)
Psychological Experiment Psychological metrics Valence entropy Automatic Lei et al. (2024)
Psychological Experiment Psychological metrics Arousal entropy Automatic Lei et al. (2024)
Psychological Experiment Performance metrics Precision of item recommendation Automatic Wang et al. (2025a)
Psychological Experiment Performance metrics Missing rate Automatic Lei et al. (2024)
Psychological Experiment Performance metrics Rejection rate Automatic Lei et al. (2024)
Psychological Experiment Internal consistency metrics Correlation between social dilemma game outcome and agent personality Automatic Bose et al. (2024)
Psychological Experiment Internal consistency metrics Behavioral similarity Automatic Li et al. (2024b)
Psychological Experiment Internal consistency metrics Perception consistency (agent perceived safety, agent perceived liveliness) LLM Verma et al. (2023)
Psychological Experiment External alignment metrics Rationality of the agent memory Automatic Wang et al. (2025a)
Psychological Experiment External alignment metrics Believability of behavior Automatic Wang et al. (2025a)
Psychological Experiment Content and textual metrics Salience of individual words Automatic De Duro et al. (2025)
Psychological Experiment Content and textual metrics Absolutist words Automatic De Duro et al. (2025)
Psychological Experiment Content and textual metrics Personal pronouns or emotions Automatic De Duro et al. (2025)
Psychological Experiment Content and textual metrics Information entropy Automatic Wang et al. (2025a)
Psychological Experiment Content and textual metrics Story (readability, personalness, redundancy, cohesiveness, likeability, believability) Human Jiang et al. (2023b)
Psychological Experiment Content and textual metrics Story (readability, personalness, redundancy, cohesiveness, likeability, believability) LLM Jiang et al. (2023b)
Simulated Individual Social and economic metrics Numbers of generated peer support strategies Automatic Liu et al. (2024b)
Simulated Individual Social and economic metrics Perceived social support questionnaire Human Liu et al. (2024b)
Simulated Individual Psychological metrics Emotions Human Pataranutaporn et al. (2024)
Simulated Individual Psychological metrics Agency Human Pataranutaporn et al. (2024)
Simulated Individual Psychological metrics Future consideration Human Pataranutaporn et al. (2024)
Simulated Individual Psychological metrics Self-reflection Human Pataranutaporn et al. (2024)
Simulated Individual Psychological metrics Insight Human Pataranutaporn et al. (2024)
Simulated Individual Psychological metrics Persona Perception Scale Human Salminen et al. (2024)
Simulated Individual Psychological metrics Persona Perception Scale Human Shin et al. (2024)
Simulated Individual Psychological metrics Persona Perception Scale Human Ha et al. (2024)
Simulated Individual Psychological metrics Persona Perception Scale Human Chen et al. (2024b)
Simulated Individual Psychological metrics Engagement Human Zhang et al. (2024a)
Simulated Individual Psychological metrics Safety Human Zhang et al. (2024a)
Simulated Individual Psychological metrics Sensitivity to personalization Automatic Giorgi et al. (2024)
Simulated Individual Psychological metrics Agent self-awareness LLM Xie et al. (2024b)
Simulated Individual Psychological metrics Personality (Big Five Invertory rated by LLM) LLM Jiang et al. (2023a)
Simulated Individual Psychological metrics Positively mention rate Automatic Kamruzzaman and Kim (2024)
Simulated Individual Psychological metrics Optimism Human Pataranutaporn et al. (2024)
Simulated Individual Psychological metrics Self-esteem Human Pataranutaporn et al. (2024)
Simulated Individual Psychological metrics Pressure perceived scale Human Liu et al. (2024b)
Simulated Individual Performance metrics Error rates (error of average, error of dispersion) Automatic Lin et al. (2024)
Simulated Individual Performance metrics Model fit indices (Chi-square to degrees of freedom ratio, Comparative Fit Index, Tucker-Lewis Index, Root Mean Square Error of Approximation) Automatic Ke and Ng (2024)
Simulated Individual Performance metrics Knowledge accuracy (WikiRoleEval with human evaluators) Human Tang et al. (2024)
Simulated Individual Performance metrics Knowledge accuracy (WikiRoleEval) LLM Tang et al. (2024)
Simulated Individual Performance metrics Win rates Automatic Chi et al. (2024)
Simulated Individual Performance metrics Comprehension Automatic Shin et al. (2024)
Simulated Individual Performance metrics Completeness Automatic Shin et al. (2024)
Simulated Individual Performance metrics Validity (average variance extracted, inter-construct correlations) Automatic Ke and Ng (2024)
Simulated Individual Performance metrics Composite reliability Automatic Ke and Ng (2024)
Simulated Individual Performance metrics Rated statement quality Human Liu et al. (2023)
Simulated Individual Performance metrics Rated statement quality LLM Liu et al. (2023)
Simulated Individual Performance metrics Conversational ability (CharacterEval) LLM Tang et al. (2024)
Simulated Individual Performance metrics Roleplay subset of MT-Bench LLM Tang et al. (2024)
Simulated Individual Performance metrics Professional scale (accuracy in replicating profession-specific knowledge) LLM Sun et al. (2024)
Simulated Individual Performance metrics Language quality LLM Zhang et al. (2024a)
Simulated Individual Performance metrics Prediction accuracy between real data and generated data (Replication success rate, Kullback-Leibler divergence) Automatic Assaf and Lynar (2024)
Simulated Individual Performance metrics Prediction accuracy between real data and generated data (Replication success rate, Kullback-Leibler divergence) Automatic Tamaki and Littvay (2024)
Simulated Individual Performance metrics Prediction accuracy between real data and generated data (Replication success rate, Kullback-Leibler divergence) Automatic Park et al. (2024)
Simulated Individual Performance metrics Prediction accuracy between real data and generated data (Replication success rate, Kullback-Leibler divergence) Automatic Yeykelis et al. (2024)
Simulated Individual Performance metrics Accuracy of distinguishing between AI-generated and human-built solutions Automatic Schuller et al. (2024)
Simulated Individual Internal consistency metrics Accuracy of reaction based on social relationship Automatic Liu et al. (2024a)
Simulated Individual Internal consistency metrics Perceived connection between personas and system outcomes Human Chen et al. (2024b)
Simulated Individual Internal consistency metrics Representativeness (Wasserstein distance, respond with similar answers to individual survey questions), Consistency (Frobenius norm, the correlation across responses to a set of questions in each survey) Automatic Moon et al. (2024)
Simulated Individual Internal consistency metrics Role consistency (WikiRoleEval with human evaluators) Human Tang et al. (2024)
Simulated Individual Internal consistency metrics Role consistency/attractiveness (WikiRoleEval, CharacterEval) LLM Tang et al. (2024)
Simulated Individual Internal consistency metrics Consistency Human Zhang et al. (2024a)
Simulated Individual Internal consistency metrics Consistency Human Mishra et al. (2023)
Simulated Individual Internal consistency metrics Future self-continuity Human Pataranutaporn et al. (2024)
Simulated Individual Internal consistency metrics Agreement between a synthetic annotator both with and without a leave-one-out attribute (Cohen’s Kappa) Automatic Castricato et al. (2024)
Simulated Individual Internal consistency metrics Consistency with the scenario and characters Automatic Zhang et al. (2024a)
Simulated Individual Internal consistency metrics Quality and logical coherence of the script content Automatic Zhang et al. (2024a)
Simulated Individual Internal consistency metrics Nation-related response percentage Automatic Kamruzzaman and Kim (2024)
Simulated Individual External alignment metrics Unknown question rejection (WikiRoleEval with human evaluators) Human Tang et al. (2024)
Simulated Individual External alignment metrics Unknown question rejection (WikiRoleEval) LLM Tang et al. (2024)
Simulated Individual External alignment metrics Accuracy of self-knowledge Automatic Liu et al. (2024a)
Simulated Individual External alignment metrics Correctness Human Zhang et al. (2024a)
Simulated Individual External alignment metrics Correctness Human Milička et al. (2024)
Simulated Individual External alignment metrics Agreement score between human raters and LLM, Automatic Liu et al. (2023)
Simulated Individual External alignment metrics Agreement score between human raters and LLM, Automatic Jiang et al. (2023a)
Simulated Individual External alignment metrics Agreement score between human raters and LLM, Automatic Liu et al. (2024a)
Simulated Individual External alignment metrics Human-likeness Human Zhang et al. (2024a)
Simulated Individual Content and textual metrics Content similarity (ROUGE-L, BERTScore, GPT-based-similarity, G-eval) Automatic Shin et al. (2024)
Simulated Individual Content and textual metrics Entity density of summarization Automatic Liu et al. (2024a)
Simulated Individual Content and textual metrics Entity recall of summarization Automatic Liu et al. (2024a)
Simulated Individual Content and textual metrics Dialog diversity Automatic Lin et al. (2024)
Simulated Individual Bias, fairness, and ethic metrics Hate speech detection accuracy Automatic Giorgi et al. (2024)
Simulated Individual Bias, fairness, and ethic metrics Population heterogeneity Automatic Murthy et al. (2024)
Simulated Society Social and economic metrics Social Conflict Count Automatic Ren et al. (2024b)
Simulated Society Social and economic metrics Social Rules Human Zhou et al. (2024b)
Simulated Society Social and economic metrics Social Rules LLM Zhou et al. (2024b)
Simulated Society Social and economic metrics Financial and Material Benefits Human Zhou et al. (2024b)
Simulated Society Social and economic metrics Financial and Material Benefits LLM Zhou et al. (2024b)
Simulated Society Social and economic metrics Converged price Automatic Toledo-Zucco et al. (2024)
Simulated Society Social and economic metrics Information diffusion Automatic Park et al. (2023)
Simulated Society Social and economic metrics Relationship formation Automatic Park et al. (2023)
Simulated Society Social and economic metrics Relationship LLM Zhou et al. (2024b)
Simulated Society Social and economic metrics Coordination within other agents Automatic Park et al. (2023)
Simulated Society Social and economic metrics Probability of social connection formation Automatic Leng and Yuan (2024)
Simulated Society Social and economic metrics Percent of social welfare maximization choices Automatic Leng and Yuan (2024)
Simulated Society Social and economic metrics Persuasion (distribution of persuasion outcomes, odds ratios) Automatic Campedelli et al. (2024)
Simulated Society Social and economic metrics Anti-social behavior (effect on toxic messages) Automatic Campedelli et al. (2024)
Simulated Society Social and economic metrics Norm Internalization Rate Automatic Ren et al. (2024b)
Simulated Society Social and economic metrics Norm Compliance Rate Automatic Ren et al. (2024b)
Simulated Society Psychological metrics NASA-TLX Scores Human Zhang et al. (2024c)
Simulated Society Psychological metrics Helpfulness rating Human Zhang et al. (2024c)
Simulated Society Psychological metrics Personality (Big Five Invertory, MBTI score, SD3 score, Linguistic Inquiry and Word Count framework, HEXACO) Automatic Frisch and Giulianelli (2024)
Simulated Society Psychological metrics Personality (Big Five Invertory, MBTI score, SD3 score, Linguistic Inquiry and Word Count framework, HEXACO) Automatic Li et al. (2024b)
Simulated Society Psychological metrics Degree of reciprocity Automatic Leng and Yuan (2024)
Simulated Society Psychological metrics Pleasure rating Human Zhang et al. (2024c)
Simulated Society Psychological metrics Trend of Favorability Decline Automatic Gu et al. (2024)
Simulated Society Psychological metrics Negative Favorability Achievement Automatic Gu et al. (2024)
Simulated Society Psychological metrics Trend of Favorability Decline Automatic Gu et al. (2024)
Simulated Society Psychological metrics Negative Favorability Achievement Automatic Gu et al. (2024)
Simulated Society Performance metrics Abstention accuracy Automatic Ashkinaze et al. (2024)
Simulated Society Performance metrics Accuracy of information gathering Automatic Kaiya et al. (2023)
Simulated Society Performance metrics Implicit reasoning accuracy Automatic Mou et al. (2024b)
Simulated Society Performance metrics Prediction accuracy (F1 score, AUC, MSE, MAE, depression risk prediction accuracy, suicide risk prediction accuracy) Automatic Lan et al. (2024)
Simulated Society Performance metrics Guess accuracy Automatic Leng and Yuan (2024)
Simulated Society Performance metrics Classification accuracy Automatic Li et al. (2024a)
Simulated Society Performance metrics Success rate Automatic Kaiya et al. (2023)
Simulated Society Performance metrics Success rate Automatic Li et al. (2023b)
Simulated Society Performance metrics Success rate Automatic Li et al. (2023b)
Simulated Society Performance metrics Success rate for coordination (identification accuracy, workflow correctness, alignment between job and agent’s skill) Automatic Li et al. (2023a)
Simulated Society Performance metrics Success rate for coordination (identification accuracy, workflow correctness, alignment between job and agent’s skill) Automatic Li et al. (2023a)
Simulated Society Performance metrics Task Accuracy Automatic Zhang et al. (2023a)
Simulated Society Performance metrics Task Accuracy Automatic Lan et al. (2024)
Simulated Society Performance metrics Errors in the prompting sequence Human Antunes et al. (2023)
Simulated Society Performance metrics Error-free execution Automatic Wang et al. (2024a)
Simulated Society Performance metrics Goal completion Human Mou et al. (2024b)
Simulated Society Performance metrics Goal completion LLM Zhou et al. (2024a)
Simulated Society Performance metrics Goal completion LLM Mou et al. (2024b)
Simulated Society Performance metrics Goal completion LLM Zhou et al. (2024b)
Simulated Society Performance metrics Efficacy Human Ashkinaze et al. (2024)
Simulated Society Performance metrics Knowledge Human Zhou et al. (2024b)
Simulated Society Performance metrics Knowledge LLM Zhou et al. (2024b)
Simulated Society Performance metrics Reasoning abilities Automatic Chen et al. (2023)
Simulated Society Performance metrics Reasoning abilities Human Chen et al. (2023)
Simulated Society Performance metrics Efficiency Automatic Piatti et al. (2024)
Simulated Society Performance metrics Text understanding and creative writing abilities (Dialogue response dataset, Commongen Challenge) LLM Chen et al. (2023)
Simulated Society Performance metrics Probabilities of receiving, storing, and retrieving the key information across the population Automatic Kaiya et al. (2023)
Simulated Society Performance metrics Correlation between predicted and real results Automatic Mitsopoulos et al. (2024)
Simulated Society Internal consistency metrics Behavioral similarity Automatic Li et al. (2024b)
Simulated Society Internal consistency metrics Semantic consistency (cosine similarity) Automatic Qiu and Lan (2024)
Simulated Society External alignment metrics Alignment (Environmental understanding and response accuracy, adherence to predefined settings) Automatic Gu et al. (2024)
Simulated Society External alignment metrics Strategy accuracy (strategies provided by the models vs. by human experts and evaluate the accuracy) Automatic Zhang et al. (2024b)
Simulated Society External alignment metrics Believability of behavior Human Zhou et al. (2024b)
Simulated Society External alignment metrics Believability of behavior Human Park et al. (2023)
Simulated Society Content and textual metrics Content similarity (ROUGE-L, BERTScore, GPT-based-similarity, G-eval, BLEU-4) Automatic Li et al. (2024a)
Simulated Society Content and textual metrics Content similarity (ROUGE-L, BERTScore, GPT-based-similarity, G-eval) Automatic Chen et al. (2024f)
Simulated Society Content and textual metrics Content similarity (ROUGE-L, BERTScore, GPT-based-similarity, G-eval) Automatic Mishra et al. (2023)
Simulated Society Content and textual metrics Semantic understanding Automatic Gu et al. (2024)
Simulated Society Content and textual metrics Complexity of generated content Automatic Antunes et al. (2023)
Simulated Society Content and textual metrics Dialogue generation quality Automatic Antunes et al. (2023)
Simulated Society Content and textual metrics Number of conversation rounds Automatic Zhang et al. (2024c)
Simulated Society Bias, fairness, and ethic metrics Bias rate (herd effect, authority effect, ban franklin effect, rumor chain effect, gambler’s fallacy, confirmation bias, halo effect) Human Liu et al. (2025)
Simulated Society Bias, fairness, and ethic metrics Bias rate (herd effect, authority effect, ban franklin effect, rumor chain effect, gambler’s fallacy, confirmation bias, halo effect) LLM Liu et al. (2025)
Simulated Society Bias, fairness, and ethic metrics Bias rate (herd effect, authority effect, ban franklin effect, rumor chain effect, gambler’s fallacy, confirmation bias, halo effect) Automatic Liu et al. (2025)
Simulated Society Bias, fairness, and ethic metrics Equality Automatic Piatti et al. (2024)
Writing Psychological metrics Qualitative feedback (expertise, social relation, valence, level of involvement) Human Benharrak et al. (2024)
Writing Performance metrics Prediction accuracy (F1 score, AUC, MSE, MAE, depression risk prediction accuracy, suicide risk prediction accuracy) Automatic Wang et al. (2024f)
Writing Performance metrics Success rate Automatic Wang et al. (2024d)
Writing Performance metrics Behavioral patterns Human Zhang et al. (2024c)
Writing Internal consistency metrics Consistency (user profile, psychotherapeutic approach) Automatic Mishra et al. (2023)
Writing Internal consistency metrics Motivational consistency LLM Wang et al. (2024d)
Writing Internal consistency metrics Audience similarity Human Choi et al. (2024)
Writing Internal consistency metrics Quality of generated dimension & values (relevance, mutual exclusiveness) Human Choi et al. (2024)
Writing External alignment metrics Factual error rate Automatic Wang et al. (2024f)
Writing External alignment metrics Correctness (politeness, interpersonal behaviour) Automatic Mishra et al. (2023)
Writing External alignment metrics Hallucination (groundedness of the chat responses) Human Choi et al. (2024)
Writing Content and textual metrics Linguistic similarity Human Choi et al. (2024)
Writing Content and textual metrics Fluency Human Mishra et al. (2023)
Writing Content and textual metrics Perplexity Automatic Mishra et al. (2023)
Writing Content and textual metrics Non-Repetitiveness Human Mishra et al. (2023)
Writing Content and textual metrics response generation quality Automatic Li et al. (2024a)
Writing Content and textual metrics Coherency LLM Wang et al. (2024d)